Highlander Cutlery manufactures kitchen knives. One of the 48 hours during a week in January. The paid an overtime rate of $18 per hour. The employees time was spent as tollows worked employee is employee earns $10 per hour for a 42-hour week. For additional hours, the Regular duties involving cutting out knife handles General shop cleanup duties 38 hours 9 hours 1 hour Idle time due to power outage
media%2F789%2F7893ad3a-638b-4772-a2b7-2b
Answer #1

Solution:

Direct labor = Regular hour worked in cutting out knife hand * regular rate= 38*10 = $380

Manufacturing overhead (Idle time) = Idle time * Regular rate = 1*10 = $10

Maufacturing overhead (Overtime premium) = Overtime hours * Over time premium = (48-42) * (18-10) = $48

Manufacturing overhead (Indirect labor) = Hours in general shop clean up * regular rate = 9*10 = $90